Evaluate the following integrals: a) ſxvx-4d c) ſtanº (4x)sec (4x) dx
Answer #1

Given as sxiz-u da let ce=8-4 = x= ut4 duzdu - scutu). Ja du. = scare thra) der = S(W² + 4531) du | 5 (facelo € 9(23)ox 2 S fby som sa de Let us for dua ih ada Es-siyang du = -1/3 sin cas du. I sin sus du = -costa) satcos (1) = cos cas COS (2+c. StanStan? (2. tancus. See canda I tancas a sechas - J(?C) -1 ). tance. See (ude Let U= seeds du z sectus tances dv= sec (u2 tanky